OpenStudy (anonymous):

Consider the following system of equations: y = −2x + 3 y = x − 5 Which description best describes the solution to the system of equations?

pooja195 (pooja195):

same thing as last time

OpenStudy (anonymous):

Here are the options: Lines y = −2x + 3 and y = 3x – 5 intersect the x-axis. Line y = −2x + 3 intersects line y = x − 5. Lines y = −2x + 3 and y = 3x − 5 intersect the y-axis. Line y = −2x + 3 intersects the origin.